### v2

# The Effects of Multi-Task Learning on ReLU Neural Network Functions

Abstract
This paper studies the properties of solutions to multi-task shallow ReLU neural network learning problems, wherein the network is trained to fit a dataset with minimal sum of squared weights. Remarkably, the solutions learned for each individual task resemble those obtained by solving a kernel method, revealing a novel connection between neural networks and kernel methods. It is known that single-task neural network training problems are equivalent to minimum norm interpolation problem in a non-Hilbertian Banach space, and that the solutions of such problems are generally non-unique. In contrast, we prove that the solutions to univariate-input, multi-task neural network interpolation problems are almost always unique, and coincide with the solution to a minimum-norm interpolation problem in a Sobolev (Reproducing Kernel) Hilbert Space. We also demonstrate a similar phenomenon in the multivariate-input case; specifically, we show that neural network learning problems with large numbers of diverse tasks are approximately equivalent to an $\ell^{2}$ (Hilbert space) minimization problem over a fixed kernel determined by the optimal neurons.
1 Introduction
This paper investigates the nature of functions learned by training neural networks on multiple “tasks” (i.e., multiple sets of labels on the same set of data points) and contrasts these solutions with those obtained by learning separate networks for each individual task. We show that the solutions of multi-task neural network learning problems can differ dramatically from the solutions to single-task problems, even if the tasks are diverse and unrelated to one another. Unlike standard intuitions Caruana (1997) and existing theory Ben-David and Schuller (2003); Maurer et al. (2016) regarding the effects and benefits of multi-task learning, our results do not rely on similarity between tasks.
We focus on shallow, vector-valued (multi-output) neural networks with Rectified Linear Unit (ReLU) activation functions, which are functions $f_{{\bm{\theta}}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{T}$ of the form
$f_{{\bm{\theta}}}({\bm{x}})\ =\ \sum_{k=1}^{K}{\bm{v}}_{k}\big{(}{\bm{w}}_{k}^{\top}{{\bm{x}}}+b_{k}\big{)}_{+}\ +\ {\bm{A}}{\bm{x}}+\bm{c}$ (1)
where $(\cdot)_{+}=\max\{0,\cdot\}$ is the ReLU activation function, ${\bm{w}}_{k}\in\mathbb{R}^{d}$, ${\bm{v}}_{k}\in\mathbb{R}^{T}$, and $b_{k}\in\mathbb{R}$ are the input and output weights and bias of the $k^{\text{th}}$ neuron. $K$ is the number of neurons and $T$ denotes the number of tasks (outputs) of the neural network. The affine term ${\bm{A}}{\bm{x}}+\bm{c}$ is the residual connection (or skip connection), where ${\bm{A}}\in\mathbb{R}^{T\times d}$ and $\bm{c}\in\mathbb{R}^{T}$. The set of all parameters is denoted by ${\bm{\theta}}:=\big{(}\{{\bm{v}}_{k},{\bm{w}}_{k},b_{k}\}_{k=1}^{K},{\bm{A}},\bm{c}\big{)}$.
Neural networks are trained to fit data using gradient descent methods and often include a form of regularization called weight decay, which penalizes the $\ell^{2}$ norm of the network weights. We consider weight decay applied only to the input and output weights of the neurons—no regularization is applied to the biases or residual connection. This is a common setting studied frequently in past work Savarese et al. (2019); Ongie et al. (2019); Parhi and Nowak (2021). Intuitively, only the input and output weights—not the biases or residual connection—affect the “regularity” of the neural network function as measured by its second (distributional) derivative, which is why it makes sense to regularize only these parameters. Given a set of training data points $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$ and a fixed width $K\geq N^{2}$, we consider the weight decay interpolation problem:  [^1] By an argument similar to the proof of Theorem 5 of Shenouda et al. (2024), the optimal objective value of (3) (hence of (2)) for any given dataset is the same for any $K\geq N^{2}$; this is the scenario that we consider in this work.
$\min_{{\bm{\theta}}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}^{2}+\|{\bm{w}}_{k}\|_{2}^{2}\ ,\ \mbox{subject to }f_{{\bm{\theta}}}({\bm{x}}_{i})={\bm{y}}_{i},\,i=1,\dots,N\ .$ (2)
By homogeneity of the ReLU activation function (meaning that $(\alpha x)_{+}=\alpha(x)_{+}$ for any $\alpha\geq 0$), the input and output weights of any ReLU neural network can be rescaled as ${\bm{w}}_{k}\mapsto{\bm{w}}_{k}/\|{\bm{w}}_{k}\|_{2}$ and ${\bm{v}}_{k}\mapsto{\bm{v}}_{k}\|{\bm{w}}_{k}\|_{2}$ without changing the function that the network represents. Using this fact, several previous works Grandvalet (1998); Grandvalet and Canu (1998); Neyshabur et al. (2015); Parhi and Nowak (2023) note that problem (2) is equivalent to
$\min_{{\bm{\theta}}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2},\ \mbox{subject to }\{\|{\bm{w}}_{k}\|_{2}=1\}_{k=1}^{K},\,f_{\bm{\theta}}({\bm{x}}_{i})={\bm{y}}_{i},\,i=1,\dots,N$ (3)
in that the minimal objective values of both training problems are the same, and any network $f_{{\bm{\theta}}}$ which solves (2) also solves (3), while any $f_{{\bm{\theta}}}$ which solves (3) also solves (2) after rescaling of the input and output weights. The regularizer $\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}$ is reminiscent of the multi-task lasso Obozinski et al. (2006). It has recently been shown to promote neuron sharing in the network, such that only a few neurons contribute to all tasks Shenouda et al. (2024).
The optimizations in (2) and (3) are non-convex and in general, they may have multiple global minimizers. As an example, consider the single-task, univariate dataset in Figure 1. For this dataset, (3) has infinitely many global solutions Savarese et al. (2019); Ergen and Pilanci (2021); Debarre et al. (2022); Hanin (2021). Two of the global minimizers are shown in Figure 1. In some scenarios, the solution on the right may be preferable to the one on the left, since the interpolation function stays closer to the training data points, and hence is more adversarially robust by most definitions Carlini et al. (2019). Moreover, recent theoretical work shows that this solution has other favorable generalization and robustness properties Joshi et al. (2024). Current training methods, however, might produce any one of the infinite number of solutions, depending on the random initialization of neural network weights as well as other possible sources of randomness in the training process. It is impossible to control this using existing training algorithms, which might explain many problems associated with current neural networks such as their sensitivity to adversarial attacks. In contrast, as we show in this paper, if we train a neural network to interpolate the data in Figure 1 along with additional interpolation tasks with different labels, then the solution is almost always unique and given by the (potentially preferable) interpolation depicted on the right. This demonstrates that multi-task learning can have a profound effect on the nature of neural network solutions.
Figure 1: Two solutions to ReLU neural network interpolation (blue) of training data (red). The functions on the left and right both interpolate the data and both are global minimizers of (2) and (3), and minimize the second-order total variation of the interpolation function Parhi and Nowak (2021). In fact, all convex combinations of the two solutions above are also solutions to this learning problem.
The main contributions of our paper are:
  1. Uniqueness of Multi-task Solutions.
    In the univariate setting ($d=1$) we prove that the solution to multi-task learning with different tasks is almost always a unique function, and we give a precise condition for the exceptional cases where solutions are non-unique.
  2. Multi-task Training $\equiv$ Kernel Method (almost always).
    In the case that the solution to the univariate weight decay problem is unique, it is given by the connect-the-dots interpolant of the training data points: i.e., the optimal solution is a linear spline which performs straight-line interpolation between consecutive data points in all tasks. On the support of the data, this solution agrees with the solution to minimum-norm interpolation in the first-order Sobolev space $H^{1}$, a reproducing kernel Hilbert space (RKHS) which contains all functions with first derivatives in $L^{2}$ De Boor and Lynch (1966). In contrast, solutions to the single-task learning problem are non-unique in general and are given by minimum-norm interpolating functions in the non-Hilbertian Banach space $\operatorname{BV}^{2}$ Parhi and Nowak (2021), which contains all functions with second distributional derivatives [^2] Technically, $\operatorname{BV}^{2}$ contains all functions with second distributional derivatives in ${\mathcal{M}}$, the space of Radon measures with finite total variation. ${\mathcal{M}}$ can be viewed as a “generalization” of $L^{1}$ (see Ch. 7.3, p.223 in Folland (1999)). in $L^{1}$. This shows that the solution to each individual task in a multi-task solution is equivalent to that of a kernel method, whereas single-task solutions generally are not.
  3. Insights on Multivariate Multi-Task Problems.
    We provide empirical evidence and mathematical analysis which indicate that similar conclusions hold in multivariate settings. Specifically, with a large number of diverse tasks, the solutions for each individual task in multi-task learning are approximately minimum-norm solutions in a particular RKHS space determined by the optimal neurons. In contrast, learning each task in isolation results in solutions that are minimum-norm with respect to a non-Hilbertian Banach norm over the optimal neurons.
2 Related Works
Function spaces associated with neural networks:
For single-output ReLU neural networks, Savarese et al. (2019); Ongie et al. (2019) related weight decay regularization on the parameters of the model to regularizing a particular semi-norm on the neural network function. Ongie et al. (2019) showed that this semi-norm is not an RKHS semi-norm, highlighting a fundamental difference between neural networks and kernel methods. Parhi and Nowak (2021, 2022); Bartolucci et al. (2023); Unser (2021) studied the function spaces associated with this semi-norm, and developed representer theorems showing that optimal solutions to the minimum-norm data fitting problem over these spaces are realized by finite-width ReLU networks. Consequently, finite-width ReLU networks trained with weight decay are optimal solutions to the regularized data-fitting problem posed over these spaces. Function spaces and representer theorems for multi-output and deep neural networks were later developed in Korolev (2022); Parhi and Nowak (2022); Shenouda et al. (2024).
Characterizations of ReLU neural network solutions:
Hanin (2021); Stewart et al. (2023) characterized the neural network solutions to (2) in the univariate input/output setting. Boursier and Flammarion (2023) showed that in the univariate input/output case, when weight decay is modified to include the biases of each neuron the solution is unique. Moreover, under certain assumptions it is often the sparsest interpolator (in terms of the number of neurons). Our work differs from these as we study the multi-task setting, showing that the solutions are almost always unique and that they are the connect-the-dots solution, which is generally not the sparsest and corresponds to minimum norm interpolation over a particular RKHS space. While characterizing solutions to (2) in the multivariate setting is more challenging, there exist some results for very particular datasets Ergen and Pilanci (2021); Ardeshir et al. (2023); zeno2024minimum.
Multi-Task Learning:
The advantages of multi-task learning has been extensively studied in the machine learning literature Obozinski et al. (2006, 2010); Argyriou et al. (2006, 2008). Theoretical properties of multi-task neural networks in particular, have also been studied in Lindsey and Lippl (2023); Collins et al. (2024); Shenouda et al. (2024). The underlying intuition in these past works has been that learning multiple related tasks simultaneously can help select or learn the most useful features for all tasks. Our work differs from this traditional paradigm as we consider multi-task neural networks trained on very general tasks which may be diverse and unrelated.
3 Univariate Multi-Task Neural Network Solutions
The “representational cost” of a function $f$ that can be represented as a neural network (1) with unit-norm input weights, i.e. $\|{\bm{w}}_{k}\|_{2}=1$ for all $k=1,\dots,K$, is defined to be
$R(f):=\inf_{\bm{\theta}}\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}\ \quad\text{s.t.}\quad f=f_{{\bm{\theta}}}$ (4)
where ${\bm{\theta}}=\left(\{{\bm{v}}_{k},{\bm{w}}_{k},b_{k}\}^{K}_{k=1},{\bm{A}},\bm{c}\right)$. Taking an inf over all possible neural network parameters is necessary as there are multiple neural networks which can represent the same function. Solutions to (3) minimize this representational cost subject to the data interpolation constraint. This section gives a precise characterization of the solutions to the multi-task neural network interpolation problem in the univariate setting ($d=1$).
For the training data points $(x_{1},{\bm{y}}_{1}),\dots,(x_{N},{\bm{y}}_{N})\in\mathbb{R}\times\mathbb{R}^{T}$, let $y_{it}$ denote the $t^{\text{th}}$ coordinate of the label vector ${\bm{y}}_{i}$. For each $t=1,\dots,T$, let ${\mathcal{D}}_{t}$ denote the univariate dataset $(x_{1},y_{it}),\dots,(x_{N},y_{Nt})\in\mathbb{R}\times\mathbb{R}$, and let
$s_{it}=\frac{y_{i+1,t}-y_{it}}{x_{i+1}-x_{i}}$ (5)
denotes the slope of the straight line between $(x_{i},y_{it})$ and $(x_{i+1},y_{i+1,t})$. The connect-the-dots interpolant of the dataset ${\mathcal{D}}_{t}$ is the function $f_{{\mathcal{D}}_{t}}$ which connects the consecutive points in dataset ${\mathcal{D}}_{t}$ with straight lines (see Figure 2). Its slopes on $(-\infty,x_{2}]$ and $[x_{N-1},\infty)$ are $s_{1t}$ and $s_{N-1,t}$, respectively.
In the following section, we state a simple necessary and sufficient condition under which the connect-the-dots interpolation $f_{{\mathcal{D}}}=(f_{{\mathcal{D}}_{1}},\dots,f_{{\mathcal{D}}_{T}})$ is the unique optimal interpolant of the datasets ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{T}$. We also demonstrate that the set of multi-task datasets which satisfy the necessary condition for non-uniqueness, viewed as a subset of $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$, has Lebesgue measure zero.
This result raises an interesting new connection between data fitting with ReLU neural networks and traditional kernel-based learning methods, because connect-the-dots interpolation is also the minimum-norm solution to a data-interpolation problem in the first-order Sobolev space $H^{1}([x_{1},x_{N}])$, itself an RKHS whose norm penalizes the $L^{2}$ norm of the function’s derivative. In particular, $f_{{\mathcal{D}}_{t}}$ agrees on $[x_{1},x_{N}]$ with the function $f(x)=\sum_{j=1}^{N}\alpha_{j}k(x,x_{j})$ whose coefficients $\alpha_{j}$ solve the kernel optimization problem
$\min_{\alpha_{1},\dots,\alpha_{N}\in\mathbb{R}}\sum_{i=1}^{N}\sum_{j=1}^{N}\alpha_{i}\alpha_{j}\,k(x_{i},x_{j})\mbox{ \ s.t. \ }\sum_{j=1}^{N}\alpha_{j}\,k(x_{i},x_{j})=y_{it},\quad i=1,\dots,N,$
with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966). Therefore, our result shows that the individual outputs of solutions to (3) for $T>1$ tasks almost always coincide on $[x_{1},x_{N}]$ with this kernel solution. In contrast, optimal solutions to the (3) in the case $T=1$ are generally non-unique and may not coincide with the connect-the-dots kernel solution Hanin (2022). We note that for $T=1$, our result is consistent with the characterization of univariate solutions to (3) in Hanin (2022).
Previous works Shenouda et al. (2024) and Lindsey and Lippl (2023) showed that multi-task training encourages neuron sharing, where all task are encouraged to utilize the same set of neurons or representations. Our result shows that univariate multi-task training is an extreme example of this phenomenon, since $f_{{\mathcal{D}}}$ can be represented using only $N-1$ neurons, all of which contribute to all of the network outputs. Therefore, in the scenario we study here, neuron sharing almost always occurs even if the tasks are unrelated.
Figure 2: The connect-the-dots interpolant $f_{{\mathcal{D}}}=(f_{{\mathcal{D}}_{1}},f_{{\mathcal{D}}_{2}},f_{{\mathcal{D}}_{3}})$ of three datasets ${\mathcal{D}}_{1},{\mathcal{D}}_{2},{\mathcal{D}}_{3}$.
3.1 Characterization and Uniqueness
Our main result is stated in the following theorem:
Theorem 3.1.
The connect-the-dots function $f_{{\mathcal{D}}}$ is always a solution to (3). Moreover, the solution to problem (3) is non-unique if and only if the following condition is satisfied: for some $i=2,\dots,N-2$, the two vectors
${\bm{s}}_{i}-{\bm{s}}_{i-1}=\frac{{\bm{y}}_{i+1}-{\bm{y}}_{i}}{x_{i+1}-x_{i}}-\frac{{\bm{y}}_{i}-{\bm{y}}_{i-1}}{x_{i}-x_{i-1}}$ (6)
and
${\bm{s}}_{i+1}-{\bm{s}}_{i}=\frac{{\bm{y}}_{i+2}-{\bm{y}}_{i+1}}{x_{i+2}-x_{i+1}}-\frac{{\bm{y}}_{i+1}-{\bm{y}}_{i}}{x_{i+1}-x_{i}}$ (7)
are both nonzero and aligned [^3] Two vectors ${\bm{u}}_{1}$ and ${\bm{u}}_{2}$ are aligned if ${\bm{u}}_{1}^{\top}{\bm{u}}_{2}=\|{\bm{u}}_{1}\|\|{\bm{u}}_{2}\|$.
If this condition not satisfied, then $f_{{\mathcal{D}}}$ is the unique solution to (3). Furthermore, as long as $T>1$ and $N>1$, the set of all possible data points $x_{1},\dots,x_{N}\in\mathbb{R}$ and ${\bm{y}}_{1},\dots,{\bm{y}}_{N}\in\mathbb{R}^{T}$ which admit non-unique solutions has Lebesgue measure zero (as a subset of $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$).
Corollary 1.
If $T>1$ and $N>1$ and the data points $x_{1},\dots,x_{N}\in\mathbb{R}$ and label vectors ${\bm{y}}_{1},\dots,{\bm{y}}_{N}\in\mathbb{R}^{T}$ are sampled from an absolutely continuous distribution with respect to the Lebesgue measure on $\mathbb{R}^{N}\times\mathbb{R}^{T\times N}$, then with probability one, the connect-the-dots function $f_{{\mathcal{D}}}$ is the unique solution to (3).
Remark 1.
The proof of Theorem 3.1, which relies mainly on Lemma 3.2 as we describe below, also characterizes solutions of the regularized loss problem
$\min_{{\bm{\theta}}}\sum_{i=1}^{N}\mathcal{L}(f_{{\bm{\theta}}}({\bm{x}}_{i}),{\bm{y}}_{i})+\lambda\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}\quad\mbox{subject to}\quad|w_{k}|=1,\>k=1,\dots,K$ (8)
for input dimension $d=1$, any $\lambda>0$, and any loss function ${\mathcal{L}}$ which is lower semicontinuous in its second argument. Specifically, any $f_{{\bm{\theta}}}$ which solves (8) is linear between consecutive data points $[x_{i},x_{i+1}]$ unless the vectors $\hat{{\bm{s}}}_{i}-\hat{{\bm{s}}}_{i-1}$ and $\hat{{\bm{s}}}_{i+1}-\hat{{\bm{s}}}_{i}$ are both nonzero and aligned, where $\hat{{\bm{s}}}_{i}:=\frac{\hat{{\bm{y}}}_{i+1}-\hat{{\bm{y}}}_{i}}{x_{i+1}-x_{i}}$ and $\hat{{\bm{y}}}_{i}:=f_{{\bm{\theta}}}(x_{i})$.
The full proof of Theorem 3.1 appears in Appendix 6.1. We outline the main ideas here. Our proof relies on the fact that any $\mathbb{R}\to\mathbb{R}^{T}$ ReLU neural network of the form (1) which solves (3) represents $T$ continuous piecewise linear (CPWL) functions, where the change in slope of the $t^{\textrm{th}}$ function at the $k^{\text{th}}$ knot is equivalent the $t^{\textrm{th}}$ entry of the $k^{\text{th}}$ output weight vector (see Appendix 6.1 for further detail). This fact allows us to draw a one-to-one correspondence between each knot in the function and each neuron in the neural network. The proof relies primarily on the following lemma
Lemma 3.2.
[…]
Then removing the knot at $x$ and instead connecting $x_{i}$ and $x_{i+1}$ by a straight line does not increase $R(f)$. Furthermore, if ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are not aligned, then doing so will strictly decrease $R(f)$.
Figure 3: The $t^{\text{th}}$ neural network output $f_{{\bm{\theta}}t}$ around the knot at $x$, where $\tau=\frac{x-x_{i}}{x_{i+1}-x_{i}}$. Each line segment in the figure is labeled with its slope. For any particular output $t$, it may be the case that $f_{{\bm{\theta}}t}$ does not have a knot at $x$ (in which case $\delta_{t}=0$); does not have a knot at $x_{i}$ (in which case $a_{t}=b_{t}+\delta_{t}$); and/or does not have a knot at $x_{i+1}$ (in which case $b_{t}-\frac{\tau}{1-\tau}\delta_{t}=c_{t}$).
Proof of Lemma 3.2.
The representational cost (4) is separable across neurons. The contribution of these neurons to $R(f)$ is:
$\|\bm{\delta}+{\bm{b}}-{\bm{a}}\|_{2}\ +\ \frac{1}{1-\tau}\|\bm{\delta}\|_{2}\ +\ \|\bm{c}-{\bm{b}}+\tau\bm{\delta}/(1-\tau)\|_{2}$
$\geq$ $\|{\bm{b}}-{\bm{a}}\|_{2}-\|\bm{\delta}\|_{2}\ +\ \frac{1}{1-\tau}\|\bm{\delta}\|_{2}\ +\ \|\bm{c}-{\bm{b}}\|_{2}-\frac{\tau}{1-\tau}\|\bm{\delta}\|_{2}$
$=$ $\|{\bm{b}}-{\bm{a}}\|_{2}\ +\ \|\bm{c}-{\bm{b}}\|_{2}$
where the inequality follows from the triangle inequality. This shows that taking $\delta_{t}=0$ for all outputs, which corresponds to connecting $x_{1}$ and $x_{2}$ with a straight line in all outputs, will never increase the representational cost of $f$. The triangle inequality used in (3.1) holds with equality for some $\bm{\delta}\neq\bm{0}$ if and only if ${\bm{a}}-{\bm{b}}$, ${\bm{b}}-\bm{c}$, and $\bm{\delta}$ are aligned with $\|\bm{\delta}\|_{2}\leq\min\{\|{\bm{a}}-{\bm{b}}\|_{2},\frac{1-\tau}{\tau}\|{\bm{b}}-\bm{c}\|_{2}\}$. ∎
Lemma 3.2 states that removing neurons which are located away from the data points and replacing them with a straight line will never increase the representational cost of the network, and it will strictly decrease the representational cost unless ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned. This result implies that the connect-the-dots interpolant $f_{{\mathcal{D}}}$ is always a solution to (3), since we may take any solution $f$ of (3) and remove all knots from it (resulting in the function $f_{{\mathcal{D}}}$) without increasing its representational cost. If ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are aligned for some $i=2,\dots,N-2$, we can view any interpolant on the interval $[x_{i},x_{i+1}]$ as an instance of Figure 3 with ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and $\bm{c}={\bm{s}}_{i+1}$. By Lemma 3.2, any neural network with a knot at some point $\tilde{x}\in(x_{i},x_{i+1})$ can have the same representational cost as the connect-the-dots solution on this interval, only if ${\bm{a}}-{\bm{b}}$ and ${\bm{b}}-\bm{c}$ are aligned.
We can also prove by contradiction that optimal solutions are unique on $[x_{i},x_{i+1}]$ as long as ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ are not aligned. Suppose that there is some other optimal interpolant $f$ which is not the connect-the-dots solution $f_{{\mathcal{D}}}$ on an interval $[x_{i},x_{i+1}]$ for which ${\bm{s}}_{i}-{\bm{s}}_{i-1}$ and ${\bm{s}}_{i+1}-{\bm{s}}_{i}$ point in different directions. Then apply the lemma repeatedly to remove all knots from $f_{\bm{\theta}}$ which are not located at the data points, except for a single remaining knot at some $\tilde{x}$ between consecutive data points. If this knot occurs after $x_{2}$ (the second data point) or before $x_{N-1}$ (second to last data point), the lemma implies automatically (again taking ${\bm{a}}={\bm{s}}_{i-1}$, ${\bm{b}}={\bm{s}}_{i}$, and ${\bm{s}}_{i+1}$) that removing this knot would strictly decrease the network’s representational cost, contradicting optimality of $f$. To conclude the proof, it remains only to show that any optimal interpolant of the dataset must agree with the connect-the-dots interpolant $f_{{\mathcal{D}}}$ before $x_{2}$ and after $x_{N-1}$; the details of this argument appear in Appendix 6.1. As our theorem and corollary quantify, real-world regression datasets (which are typically real-valued and often assumed to incorporate some random noise from an absolutely continuous distribution, e.g. Gaussian) are extremely unlikely to satisfy this special alignment condition; hence, our claim that connect-the-dots interpolation is almost always the unique solution to (3).
3.2 Numerical Illustration of Theorem 3.1
We provide numerical examples to illustrate the difference in solutions obtained from single task versus multi-task training and validate our theorem. The first row in Figure 4 shows three randomly initialized univariate neural networks trained to interpolate the five red points with minimum representational cost. While all three of the learned functions have the same representational cost (i.e., all minimize the second-order total variation subject to the interpolation constraint), they each learn different interpolants. This demonstrates that gradient descent does not induce a bias towards a particular solution. The second row shows the function learned for the first output of a multi-task neural network. This network was trained on two tasks. The first task consists of interpolating the five red points while the second consists of interpolating five randomly generated labels sampled from a standard Gaussian distribution. When trained to minimum representation cost we see that the connect-the-dots solution is the only solution learned regardless of initialization, verifying Theorem 3.1. This solution simultaneously minimizes the second-order total variation and (on the interval $[x_{1},x_{N}]$) the norm in the first-order Sobolev RKHS $H^{1}([x_{1},x_{N}])$ associated with the kernel $k(x,x^{\prime})=1-(x-x^{\prime})_{+}+(x-x_{1})_{+}+(x_{1}-x^{\prime})_{+}$ De Boor and Lynch (1966), subject to the interpolation constraints.
Figure 4: Top Row: Three randomly initialized neural networks trained to interpolate the 5 red points with minimum sum of squared weights. Bottom Row: Three randomly initialized two-output neural networks trained to interpolate a multi-task dataset with minimum sum of squared weights. The labels for the first task are the 5 red points shown while the labels for the second were randomly sampled from a standard Gaussian distribution.
4 Multivariate Multi-task Neural Network Training
In Section 3, we proved that the univariate functions learned by neural networks trained for multiple tasks can be profoundly different from the functions learned by networks trained for each task separately. In this section, we demonstrate that a similar phenomenon occurs in multivariate settings. Here we analyze neural networks of the form
$f_{{\bm{\theta}}}({\bm{x}})=\sum_{k=1}^{K}{\bm{v}}_{k}\big{(}{\bm{w}}_{k}^{\top}{\bm{x}}+b_{k}\big{)}_{+}$ (10)
where ${\bm{w}}_{k}\in{\mathbb{S}}^{d-1}$, $b_{k}\in\mathbb{R}$, ${\bm{v}}_{k}\in\mathbb{R}^{T}$, and ${\bm{\theta}}:=\{{\bm{v}}_{k},{\bm{w}}_{k},b_{k}\}_{k=1}^{K}$. Since the analysis in this section is not dependent on the residual connection, we omit it for ease of exposition. We consider the multivariate-input, $T$-task neural network training problem
$\min_{{\bm{\theta}}}\sum_{i=1}^{N}{\mathcal{L}}\left({\bm{y}}_{i},f_{{\bm{\theta}}}({\bm{x}}_{i})\right)+\lambda\sum_{k=1}^{K}\|{\bm{v}}_{k}\|_{2}$ (11)
for some dataset $({\bm{x}}_{1},{\bm{y}}_{1}),\dots,({\bm{x}}_{N},{\bm{y}}_{N})\in\mathbb{R}^{d}\times\mathbb{R}^{T}$, where ${\mathcal{L}}$ is any convex loss function which is lower semicontinuous in its second argument and separable across the $T$ tasks [^4] In other words, ${\mathcal{L}}$ is expressible as ${\mathcal{L}}({\bm{u}},{\bm{v}})=\sum_{t=1}^{T}\tilde{{\mathcal{L}}}(u_{t},v_{t})$ for any ${\bm{u}},{\bm{v}}\in\mathbb{R}^{T}$, for some univariate loss function $\tilde{{\mathcal{L}}}$. For notational convenience, we denote both the multivariate and univariate loss functions as ${\mathcal{L}}$., and $K\geq N^{2}$.
We are interested in analyzing the behavior of solutions to (11) as the number of tasks $T$ grows. If $T$ is very large and the tasks are “diverse” (meaning informally that no single task has a strong influence on any other), one might intuitively expect that each neuron of an optimal solution to (11) would contribute to many of the tasks, and thus that output weight $v_{ks}^{*}$ for an individual neuron $k$ and task $s$ would be relatively small compared to the sum of the output weights $v_{kt}^{*}$ for tasks $t\neq s$. In this case, the $k^{\textrm{th}}$ term of the regularizer in (11) would be approximately equal to
$\|{\bm{v}}_{k}^{*}\|_{2}=\sqrt{(v_{ks}^{*})^{2}+\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}}\approx\|{\bm{v}}_{k\setminus s}^{*}\|_{2}+\frac{(v_{ks}^{*})^{2}}{2\|{\bm{v}}_{k\setminus s}^{*}\|_{2}}$ (12)
for any individual task $s$, where $\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}:=\sum_{t\neq s}(v_{ks}^{*})^{2}$. The approximation above comes from the Taylor expansion $f(x)=\sqrt{x^{2}+c^{2}}=c+\frac{x^{2}}{2c}-\frac{x^{4}}{8c^{3}}+\frac{x^{6}}{16c^{5}}-\dots$, whose higher order terms quickly become negligible if $0<x\ll c$. Notice that the right hand side of (12) is a quadratic function of $v_{ks}^{*}$, which suggests that the regularization term of (11) resembles a weighted $\ell^{2}$ regularizer when $v_{ks}$ is close to its optimal value $v_{ks}^{*}$.
The above reasoning can be made precise. To formalize the notion of task diversity, we use the concept of exchangeability. A sequence $X_{1},\dots,X_{n}$ of random variables is exchangeable if their joint distribution remains unchanged for any permutation of the indices $1,\dots,n$. Intuitively, exchangeability captures the notion that the order of the random variables doesn’t matter; this is naturally consistent with most multi-task learning problems, in which the order in which the tasks are enumerated is irrelevant (it does not matter whether any particular task is labeled as task 1, task 2, etc.). Exchangeability can also be viewed as a weaker notion of independence of random variables: i.i.d. random variables are necessarily exchangeable, although the converse is not necessarily true. Another example of a random process generating exchangeable task labels is the following: let $\mu$ be a probability measure over a family of conditional probability distributions ${\cal P}$. Suppose that $p_{1},\dots,p_{T}\overset{iid}{\sim}\mu$ and ${\bm{y}}_{:t}|{\bm{x}}_{1},\dots,{\bm{x}}_{N}\sim p_{t}$, where ${\bm{y}}_{:t}:=[y_{1t},\dots,y_{Nt}]$ is the vector of labels corresponding to task $t$. Then ${\bm{y}}_{:1},\dots,{\bm{y}}_{:T}$ are exchangeable, conditioned on the data points ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$, [^5] In the remainder of the discussion, all statements are interpreted as being conditioned on ${\bm{x}}_{1},\dots,{\bm{x}}_{N}$..
Using the concept of task exchangeability, we consider the problem of minimizing
$J(v_{1s},\dots,v_{Ks}):=\sum_{i=1}^{N}{\mathcal{L}}\left(y_{is},\sum_{k=1}^{K}v_{ks}{\bm{\Phi}}_{ik}\right)+\lambda\sum_{k=1}^{K}\left\|\begin{bmatrix}v_{ks}\\ {\bm{v}}_{k\setminus s}^{*}\end{bmatrix}\right\|_{2}$ (13)
over ${\mathbb{R}}^{K}$, where $s$ is an individual task. $J$ is simply the objective function of (11) with all parameters except for $v_{1s},\dots,v_{Ks}$ held fixed at their optimal values, and with the (constant) data fitting terms from the other tasks except $s$ removed. Loosely speaking, we can view $J$ as the objective function of (11) from the perspective of a single task $s$. Note that the optimal values $v_{1s}^{*},\dots,v_{Ks}^{*}$ for (11) also minimize $J$; otherwise they would not be optimal for (11). The following theorem shows that, when the task labels are exchangeable, $J$ is approximately quadratic near its minimizer and that this approximation get stronger as the number of tasks $T$ increases.
Theorem 4.1.
Suppose that ${\bm{y}}_{:1},\dots,{\bm{y}}_{:T}$ are exchangeable and let $\mathcal{S}$ be the set of optimal active neurons solving (11), that is, the neurons for which $\|{\bm{v}}^{*}_{k}\|_{2}>0$. For an individual task $s$, consider the objective
$H(v_{1s},\dots,v_{Ks}):=\sum_{i=1}^{N}{\mathcal{L}}\left(y_{is},\sum_{k\in\mathcal{S}}v_{ks}{\bm{\Phi}}_{ik}\right)+\lambda\sum_{k\in\mathcal{S}}\left(\|\bm{v}_{k\setminus s}^{*}\|_{2}+\frac{v_{ks}^{2}}{2\|\bm{v}_{k\setminus s}^{*}\|_{2}}\right)$ (14)
where ${\bm{v}}_{k\setminus s}^{*}$ denotes the vector ${\bm{v}}_{k}^{*}$ with its $s^{\text{th}}$ element $v_{ks}^{*}$ excluded, and ${\bm{\Phi}}\in\mathbb{R}^{N\times K}$ is a matrix whose $i,k^{\text{th}}$ entry is ${\bm{\Phi}}_{ik}=\big{(}{\bm{w}}_{k}^{*}\cdot{\bm{x}}_{i}+b_{k}^{*}\big{)}_{+}$. Then with probability at least $1-O(T^{-1/3})$, the regularization term in $H$ is well-defined, and
$|J(v_{1s},\dots,v_{Ks})-H(v_{1s},\dots,v_{Ks})|\ =\ O(T^{-1/4})$ (15)
whenever $|v_{ks}|\leq T^{1/16}|v_{ks}^{*}|$ for all $k=1,\dots,K$. As a consequence, the global minimizer $v_{1s}^{\prime},\dots,v_{Ks}^{\prime}$ of $H$ satisfies
$J(v_{1s}^{\prime},\dots,v_{Ks}^{\prime})-J(v_{1s}^{*},\dots,v_{Ks}^{*})\ =\ O(T^{-1/4})$ (16)
with probability at least $1-O\left(T^{-1/3}\right)$.
The proof of Theorem 4.1 is in Appendix 6.2. The theorem states that the solution to the weighted $\ell^{2}$ minimization problem
$\min_{v_{1s},\dots,v_{Ks}}\sum_{i=1}^{N}{\mathcal{L}}\left(y_{is},\sum_{k\in\mathcal{S}}v_{ks}{\bm{\Phi}}_{ik}\right)+\frac{\lambda}{2}\sum_{k\in\mathcal{S}}\gamma_{ks}v_{ks}^{2}$ (17)
where $\gamma_{ks}:=1/\|{\bm{v}}^{*}_{k/s}\|_{2}$ is approximately optimal for the original objective (13), with stronger approximation as $T$ increases. In contrast, when $T=1$, the optimization
$\min_{v_{1},\dots,v_{K}}\sum_{i=1}^{N}{\mathcal{L}}\left(y_{i},\sum_{k=1}^{K}v_{k}\mathbf{\Psi}_{ik}\right)+\lambda\sum_{k=1}^{K}|v_{k}|.$ (18)
yields output weights which are exactly optimal for (11). Note that the matrices ${\bm{\Phi}}$ in 17 and $\mathbf{\Psi}$ in 18 are not the same, since they are determined by the optimal input weights and biases for (11), which are themselves data- and task-dependent. Nonetheless, comparing (17) and (18) highlights the different nature of solutions learned for (11) in the single-task versus multi-task case. The multi-task learning problem with exchangeable tasks favors linear combinations of the optimal neurons which have a minimal weighted $\ell^{2}$ regularization penalty. In contrast, the single-task learning problem favors linear combinations of optimal neurons which have a minimal $\ell^{1}$ penalty. Therefore, multi-task learning with a large number of diverse tasks promotes a fundamentally different linear combination of the optimal features learned in the hidden layer.
Additionally, exchangeability of the tasks allows for the following characterization of the coefficients $\gamma_{ks}$ in (17):
Lemma 4.2.
For any $s=1,\dots,T$, any $k=1,\dots,K$, and any $10<\beta<$ 1:
$\left(1-T^{\beta-1}\right)\|{\bm{v}}_{k}^{*}\|_{2}^{2}\leq\|{\bm{v}}_{k\setminus s}^{*}\|_{2}^{2}\leq\|{\bm{v}}_{k}^{*}\|_{2}^{2}$ (19)
with probability at least $1-T^{-\beta}$.
The proof is in Section 6.2. This result implies that that the weights in the weighted $\ell^{2}$ regularizer for each task converge to a common value across all the tasks with high probability as $T$ grows larger. In particular, for large $T$
$\gamma_{ks}\approx\gamma_{k}=\frac{1}{\|{\bm{v}}^{*}_{k}\|_{2}}$ (20)
with high probability. This reveals a novel connection between the problem of minimizing (13) and a norm-regularized data fitting problem in an RKHS. Specifically, consider the finite-dimensional linear space
:=${\mathcal{H}}:=\left\{f_{\bm{v}}=\sum_{k=1}^{K}v_{k}\phi_{k}:{\bm{v}}\in\mathbb{R}^{K}\right\}$ (21)
where $\phi_{k}({\bm{x}})=\left({\bm{w}}_{k}^{*}\cdot{\bm{x}}+b_{k}^{*}\right)_{+}$, equipped with the inner product
$\langle f_{\bm{v}},f_{\bm{u}}\rangle_{\mathcal{H}}={\bm{v}}^{T}{\bm{Q}}{\bm{u}}$ (22)
where ${\bm{Q}}=\textrm{diag}\left(\frac{\gamma_{1}}{2},\dots,\frac{\gamma_{K}}{2}\right)$. As a finite-dimensional inner product space, ${\mathcal{H}}$ is necessarily a Hilbert space; furthermore, finite-dimensionality of ${\mathcal{H}}$ implies that all linear functionals (including the point evaluation functional) on ${\mathcal{H}}$ are continuous. Therefore, ${\mathcal{H}}$ is an RKHS, with reproducing kernel
$\kappa({\bm{x}},{\bm{x}}^{\prime})=\sum_{k=1}^{K}\phi_{k}({\bm{x}})Q_{kk}^{-1}\,\phi_{k}({\bm{x}}^{\prime}).$ (23)
Note that $\kappa$ indeed satisfies the reproducing property, that is, $\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=f({\bm{x}})$ for any $f\in\mathcal{H}$ and any ${\bm{x}}$. To see this, write
$\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=\left\langle\sum_{k=1}^{K}Q^{-1}_{kk}\phi_{k}({\bm{x}})\phi_{k},\sum_{k=1}^{K}v_{k}\phi_{k}\right\rangle_{\mathcal{H}}$ (24)
We can view the term on the left as a function $g_{{\bm{u}}}\in\mathcal{H}$ where $u_{k}=Q_{kk}^{-1}\phi_{k}({\bm{x}})$ or ${\bm{u}}=\phi({\bm{x}}){\bm{Q}}^{-1}$, so this is equivalent to
$\langle\kappa(\cdot,{\bm{x}}),f\rangle_{\mathcal{H}}=\left\langle g_{{\bm{u}}},f\right\rangle_{\mathcal{H}}=\phi({\bm{x}}){\bm{Q}}^{-1}{\bm{Q}}{\bm{v}}=\sum_{k=1}^{K}v_{k}\phi({\bm{x}})=f({\bm{x}})$ (25)
Finding a minimizer of $H$ over $\mathbb{R}^{K}$ is thus equivalent to solving
$\operatorname*{arg\,min}_{f\in{\mathcal{H}}}\sum_{i=1}^{N}{\mathcal{L}}(y_{is},f({\bm{x}}_{i}))+\lambda\|f\|_{\mathcal{H}}^{2}.$ (26)
We provide empirical evidence for the claims presented in this section in Fig. 5 on a simple multi-variate dataset. First, we demonstrate the variety of solutions that interpolate this dataset in a single task setting. In contrast, we show that the solutions obtained via multi-task learning with additional random tasks are very similar and often much smoother than those obtained by single-task learning supporting our claim that these solutions are well approximated by a kernel method. We also verify that the optimization (14) is a good approximation for (13). We include additional experiments in Section 7 that demonstrate that these observations hold across multiple trials.
Figure 5: Top Row: Three solutions to ReLU neural network interpolation (blue surface) of training data (red). The eight data points are located at the vertices of two squares, both centered at the origin. The outer square has side-length two and values of zero at the vertices. The inner square has side-length $1$ and values of one at the vertices. All three functions interpolate the data and are global minimizers of (2) and (3) when solving for just this task (i.e., $T=1$). Bottom Row: The left figure shows the solution to the first output of a multi-task neural network with $T=101$ tasks, the first is the original task depicted in the first row while the other $100$ tasks are randomly generated from a standard Gaussian distribution. The learned solutions are almost identical across multiple optimization runs, indicating that the solutions to the multi-task optimization are nearly unique. Here we show one representative example; more examples are depicted in Appendix 7 showing that this phenomenon holds across many runs. The figure on the right shows the solution to fitting the training data by solving (14) over a fixed set of features learned by a multi-task neural network trained to fit only the $T=100$ random tasks. We observe that unlike the highly variable solutions of single-task optimization problem, the solutions obtained by solving the multi-task optimizations are nearly identical, as one would have for kernel methods. Moreover, the solution obtained by solving (14) is also similar to the solution of the full multi-task training problem with all $T=101$ tasks. Due to the simplicity of this dataset the optimality of these solutions were confirmed by solving the equivalent convex optimization to (2) developed in Ergen and Pilanci (2021).
5 Conclusion and Discussion of Limitations/Future Work
We have shown that univariate, multi-task shallow ReLU neural networks which are trained to interpolate a dataset with minimal sum of squared weights almost always represent a unique function. This function performs straight-line interpolation between consecutive data points for each task. This connect-the-dots solution is also the solution to a min-norm data-fitting problem in an RKHS. We provide mathematical analysis and numerical evidence suggesting that a similar conclusion may hold in the mulvariate-input case, as long as the tasks are sufficiently large in number and “diverse.” These results indicate that multi-task training of neural networks can produce solutions that are strikingly different from those obtained by single-task training, and highlights a novel connection between these multi-task solutions and kernel methods.
Here we study shallow ReLU networks due to their relative simplicity and amenability to mathematical analysis. Future work could aim to extend these results to other activation functions and deep network architectures. We also focus here on characterizing global solutions to the optimizations in (2) and (3). Whether or not networks trained with gradient descent-based algorithms will converge to global solutions remains an open question: our low-dimensional numerical experiments in Sections 3 and 4 indicate that they do, but a more rigorous analysis of the training dynamics would be an interesting separate line of research. Finally, while our analysis and experiments in Section 4 indicate that multivariate, multitask neural network solutions behave similarly to $\ell^{2}$ regression over a fixed kernel, we have not precisely characterized what that kernel is in the multi-input case as we have in the single-input case: developing such a characterization is of interest for future work.